OpenStudy (anonymous):

Solve the equation. |k + 6| = 3 A. {–3, 9} B. {–3, 3} C. {–9, –3} D. {all real numbers greater than or equal to –9 and less than or equal to –3}

OpenStudy (paki):

|k + 6| = 3 k + 6 = 3 and k + 6 = -3.... solve now....

OpenStudy (imstuck):

The absolute value "rules" state that the absolute value of what's inside them is equal to both the positive and negative number of the solution. If that doesn't make sense in words, this is it in number form: k + 6 = 3 or k + 6 = -3. Now solve for k and you get your two different values for k when 3 is both positive and negative. Got that?

OpenStudy (anonymous):

Yup. Thank you!